I have 4 tests and my current average is an 83%. What percentages will I have to get in order to get a 90%?
Answer by fcabanski(1391)

AVG on tests = (earned points)/(number of points possible). For 4 tests, the best possible score is 400 points (100 each).


.83 = x/400 and x = 400*.83 = 332 points.


Call the points you'll receive on the 5th test y. The total possible points in 5 tests is 500. (332+y)/500 = .9


332+y = 450


y = 118


You'd need 118 points out of 100 to achieve a 90%. Unless there's extra credit, you cannot achieve 90%.


The best possible grade you can achieve is 332+100 = 432/500 = .864 - 86.4%